40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1131     Entry time: Wed Nov 3 01:27:26 2010
Author: Koji 
Type: Laser 
Category: GYRO 
Subject: in response to the elog entries 1126-1130 

1126

- The input gain of 32.5dB=42 is good.

It seems that the amplifier starts loosing gain above 10kHz. I wonder how much phase delay we get at 10kHz.
==> We need the phase measurement too.
This may indicate that op27 is not adequately fast for our application.
A quick calculation with LISO shows the phase delay of 3deg at 10kHz.

r r1 25 gnd nm
r r2 1k nm no
op op1 op27 np nm no

uinput np 0

# uncomment this for the transfer function
#uoutput no:db:deg
# uncomment this for the noise
noisy all
noise no sum all

gnuterm pdf

freq log 100 1M 400

1127

We need the TF measurement with phase.

1128

Measurement noise floor is too high. You could use the input monitor (OP27 G=41~42) as a preamp.
The above LISO simulation showed 3.3nVrtHz. This is already pretty good preamp. 

I don't understand the dark PD noise. Is it demodulated signal? or DC noise (which is not relevant to us)?

The PD demodulated dark noise of 30nV/rtHz does not agree with our measurement last week (1uV/rtHz => 24nV/rtHz).
Probably the noise level of the FFT is affecting the measurement. USE any PREAMP

We just need the noise levels of the input monitor output in the following configuration:
- The cable terminated (Noise of the demodulation system)
- The PD connected with no light (above + noise of the PD)
- The PD connected with nominal amount of DC light - no cavity / no modulation (+shot noise)

- The PD connected with nominal amount of DC light - no cavity / with modulation (+modulation RFAM noise)

1129

The result is incomprehensible.

Use DC photocurrent (I_DC) and output noise level (V_out) for the horiz and vert axes, respectively.
For V_out, pick the noise level appropriately at an uncontaminated frequency.
Fit this curve by the model V_out = g_det sqrt(I_det + I_DC), where the parameters g_det and I_det are the equivalent transimpedance
and the equivalent noise current. g_det is a kind of gain, and I_det gives you how much DC current do you need to make the shot noise
and the system noise comperable. (i.e. You must put more light than I_det)

And the I_det is the order of the sub-mA for the low noise PD. You need to put ~10mA (= 500mW) to really test the PD shot noise.

AndI am afraid that this is not the demodulated noise.

1130

What to do

> Remeasure the cavity response (V/Hz).

No.

What we need is the precise PZT calibration and precise OPTF modeling.

Once the fast PZT is calibrated and the openloop TF is measured, we can infer the applied frequency disturbance.
Then we can easily deduce the cavity response (V/Hz).


MENU on Wednesday

1. Re-confirm the validity of the Tektronix VCO's purported external FM deviation setting.

Zach has the previous measurement but the value depends on how much output impedance the external source has.
So we test the frequency shift with the excitation we used for the IFO calibration.

2. Try to find two nice equivalent PD and characterize them.

3. Koji measures the circuit TFs while Zach tries to obtain noise plots.

4. Realignment, open loop TF measurement, CDS hooking up, etc... if possible.

 

ELOG V3.1.3-